jarielfrm139:

Polly surveyed 850 teenagers to find out their favorite type of music. She found that 32% of the teenagers surveyed like hard rock. Part A: How many teenagers surveyed like hard rock? Show your work. (4 points) Part B: In another survey, Damian found that 75% of the total number of teenagers surveyed like pop music. If 150 teenagers like pop music, what is the total number of teenagers that Damian surveyed? Show your work. (4 points) Part C: Give an example of a situation in which a percentage is greater than 100%. (2 points)

mxddi3:

ok so for part A you can do \[\frac{ x }{ 850}=\frac{ 32 }{ 100 }\] so first you do \[32\div100=t\] then take that answer and multiply it by 850. so what is 32/100 ?

mxddi3:

part B will be set up the same way, you will have \[\frac{ 150 }{ x }=\frac{ 75 }{ 100}\] then just solve by doing 100x150 then divide that answer by 75